OpenStudy (anonymous):

Solve. 2x - 5 = 13 x+18=14

OpenStudy (anonymous):

2x=13+5 can u do the rest? :)

OpenStudy (anonymous):

and x=14-18

OpenStudy (anonymous):

I don't understand how to do them though... My teacher didn't explain, nor give an example.

OpenStudy (anonymous):

ok, for 2x=13+5. add 13 and 5 then, 2x=18 divide both sides by 2 so, x=?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

correct! so now can u do the other one...

OpenStudy (anonymous):

Oops the second one is actually x+8=14. So how do I do it?

OpenStudy (anonymous):

x=14-8 x=?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

right! :)

OpenStudy (anonymous):

What about this one? 4x + 3 = 2x + 7

OpenStudy (anonymous):

4x + 3 = 2x + 7 x=2

OpenStudy (anonymous):

I am so confused I am about to cry. I am tired of this stupid algebra crap. How did you get that answer? I do not understand.

OpenStudy (anonymous):

both equal 11

OpenStudy (anonymous):

4(2)+3=11 2(2)+7=11 therefore 4(2)+3=2(2)+7

OpenStudy (anonymous):

replace x with 2

OpenStudy (anonymous):

How do I know what to replace x with for other problems though?

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

by isoclating it

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

but the answer is correct i don't have time to explain in detail sorry

OpenStudy (anonymous):

its basic math don look at it like algebra look at it like 2x2

OpenStudy (anonymous):

dont

OpenStudy (anonymous):

I am never going to get this done. Thanks for your help anyways...

OpenStudy (anonymous):

here http://www.mathway.com eneter problems for free get 5the answers but cant view the steps

OpenStudy (anonymous):

its costs extra to viw steps its free for answers

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

Thanks...